OpenStudy (anonymous):

How to I go about solving this question? given that the answer is in polar form z+3−i = 4z+9−19i, where (z+3-i) is a conjugate

Directrix (directrix):

@ineedhelpwithmatrices What are the instructions for this problem as given in your text?

OpenStudy (anonymous):

@Directrix, the instructions are "Solve for z, and give your answer in the form a+bi."

OpenStudy (anonymous):

Just solve for z like you would a normal linear equation. Get all of your z's on one side and everything else on the other z+3−i = 4z+9−19i -3z = -3 -9 + i - 19i -3z = -12 - 18i now divide both sides by -3 z = -12/-3 - 18i/-3 = 4 + 6i

OpenStudy (anonymous):

but don't you have to find the following conjugate of z+3-i? @BangkokGarrett

OpenStudy (anonymous):

Are you supposed to solve for z? If so, that's what I've done. I don't see what a conjugate has anything to do with solving for z here.

OpenStudy (anonymous):

Conjugates are usually used for dealing with imaginary numbers in the denominator of a fraction.

OpenStudy (anonymous):

well the first eqn looks like this: 

OpenStudy (anonymous):

what? you don't have ANY fractions like that in your original posting.

OpenStudy (anonymous):

You're mixed up on terminology. The CONJUGATE of 2 - i is 2 + i The CONJUGATE of 5 + 3i is 5 - 3i

OpenStudy (anonymous):

My bad

OpenStudy (anonymous):

So, I have no idea what your question is. What exactly is the equation given and the instructions?

OpenStudy (anonymous):